3-A1 – Activity Analysis

Dogwood Signs makes customized wooden signs for businesses and residences. These signs are made of wood, which the owner glues and carves by hand or with power tools. After carving the signs, she paints them or applies a natural finish. She has a good sense of her labor and materials cost behavior, but she is concerned that she does not have good measures of other support costs. Currently, she predicts support costs to be 75% of the cost materials. Close investigation of the business reveals that $70 times the number of power tool operations is a more plausible and reliable support cost relationship.

Consider estimated support costs of the following two signs that Dogwood Signs in making:

Sign A

Sign B

Materials cost

$300

$800

Number of power tool operations

10

2

?

?

  1. Prepare a report showing the support costs of both signs using each cost driver and showing the differences between the two.
  2. What advice would you give Dogwood Signs about predicting support costs?
Answer #1

Direct material cost = sign A = 300, sign B = 800

75% of such cost = sign A = 300*75/100 = 225

                            = sign B = 800*75/100 = 600

Now the support cost being calculated on the basis of $70 per tool operations

Power tool operations sign A= 10, sign B = 2

Support cost = sign A = 10* 70 = 700

                     = sign B = 2*70 = 140

The support costs are having two drivers which are completely different from each other so it is the best decision to keep drivers which help in providing the accurate results In the current scenario the direct cost is predicting a high cost to the operations that are using few power tools and low cost to the operations where many are used such that it results in losing jobs that are using few power tools and getting larger number of job tools in such operations.